f(x) = log (1+kx)/sin x ; x not equal to 0 5 ;. x=0

Grade:12

1 Answers

Vikas TU
6 years ago
Applying L-Hospital’s rule,
Differentiating numerator and denominator,
f(x)=1+kx/cosx
lim xà0 f(x)=1+0/cos 0
                   =1/1=1.